# DeRi-IGP: Learning to Manipulate Rigid Objects Using Deformable Linear Objects via Iterative Grasp-Pull

### v4

@@ Abstract
Robotic manipulation of rigid objects via deformable linear objects (DLO) such as ropes is an emerging field of research with applications in various rigid object transportation tasks. A few methods that exist in this field suffer from limited robot action and operational space, poor generalization ability, and expensive model-based development. To address these challenges, we propose a universally applicable moving primitive called Iterative Grasp-Pull (IGP). We also introduce a novel vision-based neural policy that learns to parameterize the IGP primitive to manipulate DLO and transport their attached rigid objects to the desired goal locations. Additionally, our decentralized algorithm design allows collaboration among multiple agents to manipulate rigid objects using DLO. We evaluated the effectiveness of our approach in both simulated and real-world environments for a variety of soft-rigid body manipulation tasks. In the real world, we also demonstrate the effectiveness of our decentralized approach through human-robot collaborative transportation of rigid objects to given goal locations. We also showcase the large operational space of IGP primitive by solving distant object acquisition tasks. Lastly, we compared our approach with several model-based and learning-based baseline methods. The results indicate that our method surpasses other approaches by a significant margin. The project supplementary material and videos are available at: https://sites.google.com/view/deri-igp/home
@@ III-C Action Generation Module
Action Sampling Strategy. Given the predicted pulling point $\eta_{t}^{i}$ of the IGP action $\phi_{t}^{i}=(\lambda_{t}^{i},~{}\eta_{t}^{i})$, we sample $m$ extra IGP actions $\hat{\Phi}_{t}^{i}=\{(\lambda_{t}^{i},~{}\eta_{t}^{i:1}),~{}(\lambda_{t}^{i},~{}\eta_{t}^{i:2}),~{}\dots,~{}(\lambda_{t}^{i},~{}\eta_{t}^{i:m})\}$ from a Gaussian distribution centered at ${}\eta_{t}^{i}$ with a predefined variance 0.5 $meter^{2}$, i.e., $\hat{\Phi}_{t}^{i}\sim\mathcal{N}(\phi_{t}^{i},~{}0.5)$.
@@ III-F Multiagent Workflow
  • GPN and PPN collaboratively propose an IGP action for each of the $K$ agents.
  • The Gaussian sampler generates $m$ additional commands around each proposed action, leading to a total of $K\times(m+1)$ samples.
@@ III-H Data Generation
We use MuJoCo [25] to build our simulated setup. The data generation process of DeRi-IGP is similar to [1]. However, we employ a single-robot setup to accommodate the extensive operational space. Additionally, we limit the max velocity of the robot end-effector to avoid sliding and drifting. It is important to emphasize that in this study, we deliberately constrain the action speed of agents to uphold a quasi-static process. Consequently, acceleration in the process is negligible since the moving speed of the object is slow. Our setup necessitates frequent resets of the data collection environment, placing the rigid object at random distant positions and randomizing the environment configurations. For each move, we sample and execute $\phi$ and collect the initial and final depth map, the color map, and the rigid object’s positions. After processing, we form an instance of the dataset, i.e., ($\phi_{t}$, $d_{t}$, $s_{t}$, $r_{t}$, $v_{i}$, $p_{i}$). We collected 20,000 instances and split them into training, validation, and testing with 0.8: 0.1: 0.1 ratios.
@@ IV Experiments
This section presents our evaluation analysis for DeRi-IGP in solving goal-reaching and long-distance acquisition tasks. In all experiments, we vary the rope length and object dimensions randomly.In both the simulator and real world, we employ realsense D455 cameras and UR5e arms with its built-in positional controller to conduct the experiments.
@@ IV-C1 Simulated goal-reaching task evaluation
Table I summarizes the results, accompanied by a visual depiction of a task in Fig. 2. It can be seen that the other methods outperform DeRi-IGP-NS by a considerable margin, providing a compelling rationale for the use of the DPN-based action sampling framework. The step cost of DeRi-IGP-NS is low because it solved fewer cases than other methods, mostly with a short horizon. Regarding grasping point selection, the relatively low performance of DeRi-IGP-RG indicates the importance of GPN. It also reflects the efficiency of PPN. Specifically, even with a random grasping point, PPN can still generate a reasonable pulling point to finish the task. Furthermore, by evaluating subgoal planners, the results suggest comparable performance levels among them. However, GIP notably takes fewer steps than LGP and demonstrates higher stability, as indicated by FO. In addition, we exclude PPN’s ablation since in DeRi-Bot [1], the informed sampling and random sampling baseline for pulling action have already been shown to exhibit poorer performance than the data-driven neural network-based pulling approach. Finally, the two model-based baselines, TDA and G-TDA, exhibit relatively poor performance than our method. It is important to highlight that these model-based baselines directly utilize simulator feedback to determine both the pulling distance and direction. In contrast, all other DeRi-IGP baselines rely on relatively low-resolution depth images for spatial information. This outcome indicates that tension-driven, model-based methods, despite having accurate system information, struggle to solve the given task. Note that we exclude DeRi-Bot [1] as it lacks grasping primitive and cannot generalize to varying rope lengths and environment size and setup.
@@ IV-C3 Real-world Human-Robot Evaluation
We conduct the real-world goal-reaching task under the human-robot operational configuration [^1] The human-robot collaboration study in this paper has been approved by the Purdue Institutional Review Board (Purdue IRB-2024-1299). We follow the related protocols to ensure the human subjects’ privacy, confidentiality, and anonymity.. An example of such a task is presented in Fig. 3. Such a task assesses the sim-to-real transfer capability and the human-robot team compatibility of DeRi-IGP. Note that DeRi-IGP exhibits an asynchronous behavior, so it is expected to naturally collaborate with human operators. This experiment follows most procedures used in previous inter-robot tasks, with the main difference being that the human operator can choose to manipulate the rope or skip and let the robot act again after it finishes an action.

### v5

@@ Abstract
Robotic manipulation of rigid objects via deformable linear objects (DLO) such as ropes is an emerging field of research with applications in various rigid object transportation tasks. A few methods that exist in this field suffer from limited robot action and operational space, poor generalization ability, and expensive model-based development. To address these challenges, we propose a universally applicable moving primitive called Iterative Grasp-Pull (IGP). We also introduce a novel vision-based neural policy that learns to parameterize the IGP primitive to manipulate DLO and transport their attached rigid objects to the desired goal locations. Additionally, our decentralized algorithm design allows collaboration among multiple agents to manipulate rigid objects using DLO. We evaluated the effectiveness of our approach in both simulated and real-world environments for a variety of soft-rigid body manipulation tasks. In the real world, we also demonstrate the effectiveness of our decentralized approach through human-robot collaborative transportation of rigid objects to given goal locations. We also showcase the large operational space of IGP primitive by solving distant object acquisition tasks. Lastly, we compared our approach with several model-based and learning-based baseline methods. The results indicate that our method surpasses other approaches by a significant margin. The project supplementary material is available at: https://sites.google.com/view/deri-igp
Index Terms:
Deep Learning in Grasping and Manipulation, Imitation Learning, Human-Robot Collaboration
@@ III-C Action Generation Module
Action Sampling Strategy. Soft body dynamics exhibit high complexity and sensitivity to perturbations. A minor change in action can lead to significantly divergent outcomes. Our model predictive scheme addresses this uncertainty by evaluating multiple actions distributed around the neural network’s initial prediction. It provides a probabilistic approach to action selection, improving robustness beyond a single network prediction. Specifically, given the predicted pulling point $\eta_{t}^{i}$ of the IGP action $\phi_{t}^{i}=(\lambda_{t}^{i},~{}\eta_{t}^{i})$, we sample $m$ extra IGP actions $\hat{\Phi}_{t}^{i}=\{(\lambda_{t}^{i},~{}\eta_{t}^{i:1}),~{}(\lambda_{t}^{i},~{}\eta_{t}^{i:2}),~{}\dots,~{}(\lambda_{t}^{i},~{}\eta_{t}^{i:m})\}$ from a Gaussian distribution centered at ${}\eta_{t}^{i}$ with a predefined variance 0.5 $meter^{2}$, i.e., $\hat{\Phi}_{t}^{i}\sim\mathcal{N}(\phi_{t}^{i},~{}0.5)$.
@@ III-F Multiagent Workflow
  • GPN and PPN collaboratively propose an IGP action for each of the $n$ agents.
  • The Gaussian sampler generates $m$ additional commands around each proposed action, leading to a total of $n\times(m+1)$ samples.
@@ III-H Data Generation
We use MuJoCo [25] to build our simulated setup. The data generation process of DeRi-IGP is similar to [1]. However, we employ a single-robot setup to accommodate the extensive operational space. Additionally, we limit the max velocity of the robot end-effector to avoid sliding and drifting. It is important to emphasize that in this study, we deliberately constrain the action speed of agents to uphold a quasi-static process. Consequently, acceleration in the process is negligible since the moving speed of the object is slow. Our setup necessitates frequent resets of the data collection environment, placing the rigid object at random distant positions and randomizing the environment configurations. We randomize the dimension of the rigid object for generalization. For each move, we sample and execute $\phi$ and collect the initial and final depth map, the color map, and the rigid object’s positions. After processing, we form an instance of the dataset, i.e., ($\phi_{t}$, $d_{t}$, $s_{t}$, $r_{t}$, $v_{i}$, $p_{i}$). We collected 20,000 instances and split them into training, validation, and testing with 0.8: 0.1: 0.1 ratios.
@@ IV Experiments
This section presents our evaluation analysis for DeRi-IGP in solving long-distance and goal-reaching acquisition tasks. In all experiments, we vary the rope length and object dimensions randomly. The box dimensions are sampled uniformly within the following ranges, measured in meters: the depth ranges from 0.10 to 0.15, the height ranges from 0.05 to 0.10, and the width ranges from 0.10 to 0.20. The length of the rope is uniformly sampled between 2.5 and 4.0 meters. In both the simulator and real world, we employ realsense D455 cameras and UR5e arms with its built-in positional controller to conduct the experiments.
@@ IV-A Comparison Methods
Note, we set $m=10$ for all the methods with the action sampling process.
@@ IV-C1 Simulated goal-reaching task evaluation
Table I summarizes the results, accompanied by a visual depiction of a task in Fig. 2. It can be seen that the other methods outperform DeRi-IGP-NS by a considerable margin, providing a compelling rationale for the use of the DPN-based action sampling framework. The step cost of DeRi-IGP-NS is low because it solved fewer cases than other methods, mostly with a short horizon. Regarding grasping point selection, the relatively low performance of DeRi-IGP-RG indicates the importance of GPN. It also reflects the efficiency of PPN. Specifically, even with a random grasping point, PPN can still generate a reasonable pulling point to finish the task. In addition, DeRi-IGP-RG outperforms DeRi-IGP-NS, likely due to differences in action space complexity. While pulling points cover a broader range, nearly equivalent to the robot arm’s operational space, grasping points are more constrained along the rope, which explains the varying performance and sensitivity of pulling and grasping point selections. Furthermore, by evaluating subgoal planners, the results suggest comparable performance levels among them. However, GIP notably takes fewer steps than LGP and demonstrates higher stability, as indicated by FO. In addition, we exclude PPN’s ablation since in DeRi-Bot [1], the informed sampling and random sampling baseline for pulling action have already been shown to exhibit poorer performance than the data-driven neural network-based pulling approach. Finally, the two model-based baselines, TDA and G-TDA, exhibit relatively poor performance than our method. It is important to highlight that these model-based baselines directly utilize simulator feedback to determine both the pulling distance and direction. In contrast, all other DeRi-IGP baselines rely on relatively low-resolution depth images for spatial information. This outcome indicates that tension-driven, model-based methods, despite having accurate system information, struggle to solve the given task. Note that we exclude DeRi-Bot [1] as it lacks grasping primitive and cannot generalize to varying rope lengths and environment size and setup.
@@ IV-C3 Real-world Human-Robot Evaluation
We conduct the real-world goal-reaching task under the human-robot operational configuration [^1] The human-robot collaboration study in this paper has been approved by the Purdue Institutional Review Board (Purdue IRB-2024-1299). We follow the related protocols to ensure the human subjects’ privacy, confidentiality, and anonymity.. An example of such a task is presented in Fig. 3. Such a task assesses the sim-to-real transfer capability and the human-robot team compatibility of DeRi-IGP. Note that DeRi-IGP exhibits an asynchronous behavior, so it is expected to naturally collaborate with human operators. In our experiment, the human operator holds the highest priority during task execution. The robot will only take action if the human operator does not take action within a specified time threshold.